OpenStudy (anonymous):

What is the limiting reactant when 4.5 moles of aluminum react with 6.7 moles of oxygen gas? Unbalanced equation: Al + O2 â ' Al2O3 Show, or explain, all of your work along with the final answer.

OpenStudy (anonymous):

The things in place of the invalid character symbols were a ` and an a with an accent mark.

OpenStudy (anonymous):

well if the reaction product is Al2O3, then 2 moles of aluminium will react with 3 moles of oxygen atoms

OpenStudy (anonymous):

So the limiting reactant would be oxygen?

OpenStudy (anonymous):

you've got 6.7 moles of oxygen gas, how many moles of oxygen atoms is that ?

OpenStudy (anonymous):

3?

OpenStudy (anonymous):

??????????????????????? how many oxygen atoms make up an oxygen molecule ?

OpenStudy (anonymous):

Oh, two.

OpenStudy (anonymous):

right, so however many moles of oxygen molecules you've got, there are twice as many moles of oxygen atoms

OpenStudy (anonymous):

and for every 3 moles of oxygen atoms, you're going to need 2 moles of aluminium

OpenStudy (michele_laino):

The balanced reaction is: \[2Al+\frac{ 3 }{ 2 }O _{2}=Al _{2}O _{3}\], or, dividing by two: \[Al+\frac{ 3 }{ 4 }O _{2}=\frac{ 1 }{ 2 }Al _{2}O _{3}\] From the last equation, we have that 1 mole of Al, reacts, with 3/4 moles of O_2, in order to get 1/2 mole of Al_2O_3, so 4.5 moles of Al, are able to reacts only with 3/4*4.5=3.4 moles of O_2, in order to get 1/2*4.5=2.25 moles of Al_2O_3. FInally the excess of O_2, is 6.7-3.4=3.3 moles of O_2